OpenStudy (i_ask_alot):

How many minutes is 1.5 hours? I medal and fan!!!

OpenStudy (jdoe0001):

well, how many minutes in 1hr?

OpenStudy (samanthagreer):

do you know haw many minutes is in 1 hour

OpenStudy (i_ask_alot):

@jdoe0001 @samanthagreer 60

OpenStudy (samanthagreer):

ok so whaat is half of 60

OpenStudy (i_ask_alot):

@samanthagreer 30

OpenStudy (samanthagreer):

ok soadd 60 and 30 together

OpenStudy (jdoe0001):

1hr = 60 1/2 hr = 60/2 1 + 1/2 hr = 1.5 hr = 60 + 60/2

OpenStudy (thecalchater):

60 minutes in 1 hour...... 1/2 hour times 60 minutes is 30 60+30 is 90 minutes

OpenStudy (samanthagreer):

add 60 and 30

OpenStudy (i_ask_alot):

I have fanned and gave you a medal @I_love_my_nieces I will fan you all

OpenStudy (samanthagreer):

@I_love_my_nieces please dont give direct anwsers

OpenStudy (samanthagreer):

@I_Ask_Alot what do you think it is what is 60 +30 =

OpenStudy (123ab456c):

sup adriannna

OpenStudy (i_ask_alot):

@samanthagreer 90

OpenStudy (samanthagreer):

yes great job so that is your anwser

OpenStudy (i_ask_alot):

Thanks

OpenStudy (samanthagreer):

no proablem any time

OpenStudy (123ab456c):

hey adrianna its alex!:):)

OpenStudy (samanthagreer):

please dm the user if you with to comuivat with then please dont do it on a post were the user is trying to get help

OpenStudy (123ab456c):

uhhhh excusme who do u think ur talking to

OpenStudy (tootzrll):

please try to be respectful to other openstudiers, she was only trying to help thanks :) @123AB456C

OpenStudy (123ab456c):

ehhhhhhhhhhhhh